# Subgraph-Aware Training of Text-based Methods for Knowledge Graph Completion

### v3

@@ 4.2 Subgraph as a Mini-Batch
In this subsection, we describe our training framework dubbed Subgraph as a Mini-batch (SaaM). We count the number of visits for all triples in the training set $T$ throughout the training process. For every iteration, we select the subgraph whose center triple is the least frequently visited, to prioritize unpopular and peripheral triples. The rationale behind this selection will be elaborated in Section 6.2. Next, we randomly select $|B|/2$ triples from the subgraph, and feed to the model a mini-batch $B$ of the these selected triples $(h,r,t)$ and their inverse triples $(t,r^{-1},h)$. For every positive triple $(h,r,t)\in B$, we obtain negative triples $(h,r,\hat{t})$ with $t$ replaced by $|B|-1$ tails $\hat{t}$ of the other triples in $B$. As per our observation in Figure 1, these negative triples are likely to be hard negatives, which will facilitate contrastive learning.
To introduce a concept corresponding to an epoch in our training framework, let iterating above process for $|T|/|B|$ times, i.e., selecting and feeding $|T|/|B|$ subgraphs, be a “phase” as opposed to an epoch that visits every triple exactly once.
4.3 Subgraph-Aware Contrastive Learning
For effective contrastive learning, we propose to incorporate two structure-aware factors into InfoNCE loss with additive margin (Chen et al., 2020; Yang et al., 2019) over a mini-batch: (a) structural hardness of negatives for each positive triple, and (b) structural hardness of head entities in the mini-batch. These two factors are simply obtained from the topological structure of the subgraph, while incurring a minimal computational overhead overall.
4.3.1 Structural Hardness of Negative Triples
@@
Entities close to each other are more likely to be related than entities far away from each other. Although text-based KGC methods capture semantic relations within the text of triples, they overlook capturing the proximity between two entities in a negative triple in KGs. To tackle this problem, we propose a novel InfoNCE loss that penalizes structurally hard negative triples. For each positive triple $(h,r,t)$ in a mini-batch $B$ of SaaM, the loss function $L_{(h,r,t)}$ is defined as
$L_{(h,r,t)}=-\log\frac{e^{(\phi(h,r,t)-\gamma)/\tau}}{e^{(\phi(h,r,t)-\gamma)/\tau}+\sum\limits_{i=1}^{|B|-1}e^{(\phi(h,r,t_{i})+\beta\omega_{ht_{i}})/\tau}}\end{aligned}$
@@
where $\gamma$ is an additive margin, a temperature parameter $\tau$ adjusts the importance of negatives, structural hardness $\omega_{ht_{i}}$ stands for how hard a negative triple ($h,r,t_{i}$) is in terms of the structural relation between $h$ and $t_{i}$ in the KG, and $\beta$ is a trainable parameter that adjusts the relative importance of $\omega_{ht_{i}}$. Specifically, we define $\omega_{ht_{i}}$ as the reciprocal of the distance (i.e., length of the shortest path) between $h$ and $t_{i}$ to impose a larger weight $\omega_{ht_{i}}$ to the negative triple with a shorter distance between $h$ and $t_{i}$, which serves as a harder negative triple.
@@
Since computing the exact distance between every head and every tail in $B$ may spend considerable time, in our implementation, we calculate the approximate distance between $h$ and $t_{i}$, i.e., the multiplication of two distances: (d1) the distance between $h$ and head $h_{c}$ of the center triple of $B$, and (d2) the distance between $t$ and $h_{c}$. For this, the distance between $h_{c}$ and every entity in $B$ is pre-computed before training, the multiplication between the two distances is performed in parallel during training, which requires a minimal computational overhead. [^4] We conducted experiments using the exact distance and different approximate distances, e.g., the sum of (d1) and (d2), but the performance gap between all the methods is very marginal.
4.3.2 Structural Hardness of Entities
@@
To encourage the text-based method to more sensitively adapt to varying relations for many triples with the identical head, we propose a novel loss weighting strategy that penalizes structurally hard head entities. For each triple in a mini-batch $B$ in SaaM, mini-batch loss $L_{B}$ is defined as
$L_{B}=\sum_{(h,r,t)\in B}\psi_{h}L_{(h,r,t)}$
where structural hardness $\psi_{h}$ of head $h$ indicates how difficult $(h,r,t)$ is in terms of the structural property of $h$ in the KG. Specifically, we define $\psi_{h}$ as $log(d_{h}+1)$ where $d_{h}$ represents the degree of $h$, and adding one to $d_{h}$ can prevent $\psi_{h}$ from becoming zero when $d_{h}=1$. To sum up, $\psi_{h}$ ensures that the triples whose heads have a larger degree contribute more significantly to $L_{B}$.
@@
Table 1: Statistics of datasets.
Approach WN18RR NELL-995 FB15k-237
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
@@
TuckER Balazevic et al. (2019) 0.466 0.432 0.478 0.518 0.423 0.363 0.455 0.536 0.361 0.265 0.391 0.538
RotatE Sun et al. (2019) 0.471 0.421 0.490 0.568 0.411 0.350 0.439 0.522 0.335 0.243 0.374 0.529
KGTuner Zhang et al. (2022) 0.481 0.438 0.499 0.556 0.428 0.371 0.458 0.544 0.345 0.252 0.381 0.534
CompoundE Ge et al. (2023) 0.492 0.452 0.510 0.570 0.434 0.381 0.466 0.550 0.350 0.262 0.390 0.547
CSProm-KG Chen et al. (2023) 0.569 0.520 0.590 0.675 0.422 0.381 0.474 0.535 0.355 0.261 0.389 0.531
@@
StAR Wang et al. (2021a) 0.398 0.238 0.487 0.698 0.419 0.238 0.395 0.433 0.288 0.195 0.313 0.480
HaSa Zhang et al. (2023a) 0.538 0.444 0.588 0.713 0.411 0.310 0.431 0.524 0.304 0.220 0.325 0.483
KG-S2S Chen et al. (2022) 0.572 0.529 0.595 0.663 0.416 0.311 0.421 0.519 0.337 0.255 0.374 0.496
SimKGC Wang et al. (2022a) 0.671 0.580 0.729 0.811 0.425 0.353 0.447 0.531 0.340 0.252 0.365 0.515
GHN Qiao et al. (2023) 0.678 0.596 0.719 0.821 - - - - 0.339 0.251 0.364 0.518
Ensemble approach
StAR Wang et al. (2021a) 0.520 0.456 0.509 0.707 0.415 0.311 0.402 0.510 0.332 0.229 0.387 0.526
SATKGC-SPW-DW 0.672 0.605 0.714 0.811 0.411 0.310 0.431 0.524 0.345 0.225 0.367 0.520
SATKGC-DW 0.679 0.609 0.723 0.813 0.429 0.361 0.466 0.551 0.351 0.261 0.372 0.525
SATKGC-SPW 0.681 0.612 0.729 0.820 0.431 0.369 0.472 0.556 0.359 0.271 0.393 0.538
SATKGC 0.689 0.621 0.731 0.823 0.433 0.389 0.476 0.560 0.368 0.276 0.401 0.548
Table 2: KGC results for the WN18RR, NELL-995, and FB15k-237 datasets. “SPW” and “DW” refer to shortest path weight and degree weight respectively. The best and second-best performances are denoted in bold and underlined respectively. $\dagger$: numbers are from Qiao et al. (2023).
@@
TransE Bordes et al. (2013) 0.253 0.170 0.311 0.392 - - - -
RotatE Sun et al. (2019) 0.290 0.234 0.322 0.390 - - - -
@@
DKRL Xie et al. (2016) 0.160 0.120 0.181 0.229 0.231 0.059 0.320 0.546
KEPLER Wang et al. (2021c) 0.212 0.175 0.221 0.276 0.403 0.225 0.517 0.725
BLP-ComplEx Daza et al. (2021) - - - - 0.491 0.261 0.670 0.881
BLP-SimplE Daza et al. (2021) - - - - 0.490 0.283 0.641 0.868
SimKGC Wang et al. (2022a) 0.358 0.313 0.376 0.441 0.714 0.609 0.785 0.917
SATKGC 0.404 0.361 0.423 0.479 0.763 0.659 0.809 0.927
Table 3: KGC results for Wikidata5M-Trans (transductive setting) and Wikidata5M-Ind (inductive setting). The results for embedding-based approach on Wikidata5M-ind are missing as they cannot be used in the inductive setting. Additionally, BLP-ComplEx Daza et al. (2021) and BLP-SimplE Daza et al. (2021) results on Wikidata5M-Trans are missing because they are inherently targeted for inductive KGC.
5 Experiments
5.1 Experimental Setup
@@
KoPA (Zhang et al., 2023c) 0.483 0.344 0.559 0.721
SATKGC 0.668 0.557 0.742 0.858
@@
Table 4: KGC results on FB15K-237N. A correct tail is ranked among 1,000 randomly-selected entities due to the long inference time of KoPA.
For evaluation we adopt widely-used KG datasets WN18RR, FB15k-237, NELL-995 and Wikidata5M. Table 1 shows their statistics, and more details are provided in Appendix C.
For every incomplete triple in the test set, we compute mean reciprocal rank (MRR) and Hits@$k$ where $k\in\{1,3,10\}$ as evaluation metrics based on the rank of the correct entity among all the entities in the KG. We use the mean of the forward and backward prediction results as the final performance measure. Implementation details are described in Appendix D.
5.2 Main Results
We compare SATKGC with existing embedding-based and text-based approaches. Table 2 shows the results on WN18RR, NELL-995, and FB15k-237, and Table 3 shows the results on Wikidata5M-Trans and Wikidata5M-Ind. [^5] The results of GHN on NELL-995 are unavailable due to a lack of access to the requisite codes. SATKGC denotes the bi-encoder architectures trained by our learning framework in Figure 3. Note that Table 2 presents the performance of the latest baselines, and the complete results for all baselines compared with ours are found in Appendix K. SATKGC consistently outperforms all the existing methods on all the datasets except MRR on NELL-995. SATKGC demonstrates significantly higher MRR and Hits@1 than other baselines especially on WN18RR and FB15k-237. The small performance improvement of our method on NELL-995 stems from the limited textual context in this dataset, unlike other datasets. Nevertheless, SATKGC is the runner-up lagged only 0.001 behind CompoundE for MRR on NELL-995, while outperforming all baselines on the remaining metrics.
As shown in Table 3, SATKGC demonstrates its applicability to large-scale KGs, and achieves strong performance in both inductive and transductive settings. [^6] Baselines listed in Table 2 but not in Table 3 could not be evaluated on Wikidata5M due to out-of-memory for StAR and CSProm-KG, or unreasonably large training time, i.e., more than 100 hours expected, for the remaining baselines. SATKGC significantly outperforms the baselines on Wikidata5M-Ind, reaching MRR of 0.763 and Hits@1 of 0.659.777Wikidata5M-Ind shows better performance than Wikidata5M-Trans, because a model ranks 7,475 entities in the test set for Wikidata5M-Ind while ranking about 4.6 million entities for Wikidata5M-Trans. In the transductive setting, performance degrades in the order of WN18RR, Wikidata5M-Trans, and FB15k-237, showing that a higher average degree of entities tends to negatively affect performance.
@@ 5.2 Main Results
To compare our framework employing BERTs with a LLM-based model, we evaluate the KGC performance of KoPA (Zhang et al., 2023c), which adopts Alpaca (Taori et al., 2023) fine-tuned with LoRA (Hu et al., 2021) as its backbone, on the FB15k-237N dataset (Zhang et al., 2023c). Since KoPA cannot perform inference for all queries within a reasonable time, i.e., 111 hours expected, we rank the correct tail among 1,000 randomly selected entities for both KoPA and SATKGC. As shown in Table 4, SATKGC outperforms KoPA on all metrics, demonstrating that LLMs do not necessarily produce superior results on KGC.
5.3 Ablation study
To show the contributions of applying SaaM and adding the two weighting factors to InfoNCE loss, we compare results of four different settings including SATKGC in Table 2: SATKGC-SPW-DW denotes applying only SaaM and using original InfoNCE loss, SATKGC-SPW represents applying SaaM and the degree factor $\psi_{h}$, and SATKGC-DW represents applying SaaM and the distance factor $\omega_{ht_{i}}$. SATKGC-SPW-DW already shows higher Hits@1 than other baselines on WN18RR, which highlights that SaaM alone can lead to performance improvement. The efficacy of SaaM is further evidenced in Appendix F. Between SATKGC-SPW and SATKGC-DW, SATKGC-SPW achieves higher performance, which indicates that the degree factor contributes more than the distance factor.

### v5

@@ 4.2. Subgraph as a Mini-batch
We now present a way to select one of the sampled subgraphs, and utilize that subgraph as a mini-batch, dubbed Subgraph as a Mini-batch (SaaM). For every iteration, we select the subgraph whose center triple has been the least frequently visited, to prioritize unpopular and peripheral triples. For this, we count the number of visits for all triples in the training set $\mathcal{T}$ throughout the training process. The rationale behind this selection will be elaborated in Section 5.5.
Next, we randomly select $|\mathcal{B}|/2$ triples from the subgraph, and then feed to the bi-encoders a mini-batch $\mathcal{B}$ consisting of the selected triples $(h,r,t)$ and their corresponding inverse triples $(t,r^{-1},h)$. For every positive triple $(h,r,t)\in\mathcal{B}$, we obtain negative triples $(h,r,\hat{t})$ with $t$ replaced by $|\mathcal{B}|-1$ tails $\hat{t}$ of the other triples in $\mathcal{B}$. As per our observation in Figure 1, these negative triples are likely to be hard negatives, which will facilitate contrastive learning. As a result, we end up with iterating the above process, i.e., selecting the subgraph and feeding the triples in that subgraph, $|\mathcal{T}|/|\mathcal{B}|$ times for each epoch.
4.3. Proximity-aware Contrastive Learning
@@
Most PLMs for KGC overlook capturing the proximity between two entities in a negative triple in the KG, though they capture semantic relations within the text of triples, as described in the first observation of Section 1. For effective contrastive learning for these methods, we incorporate a topological characteristic, i.e., the proximity between two entities, of the KG into InfoNCE loss with additive margin (Chen et al., 2020; Yang et al., 2019) by measuring how hard a negative triple is in the KG. For each positive triple $(h,r,t)$ in a mini-batch $\mathcal{B}$, we propose loss $\mathcal{L}_{(h,r,t)}$ below based on our aforementioned observation, i.e., entities close to each other are more likely to be related than entities far away from each other:
$\mathcal{L}_{(h,r,t)}=-\log\frac{\text{exp}(\frac{\phi_{\mathcal{G}}(h,r,t)-\gamma}{\tau})}{\text{exp}(\frac{\phi_{\mathcal{G}}(h,r,t)-\gamma}{\tau})+\sum\limits_{i=1}^{|B|-1}\text{exp}(\frac{\phi_{\mathcal{G}}(h,r,t_{i})}{\tau})}\end{aligned}$
$\phi_{\mathcal{G}}(h,r,t_{i})=\text{cos}(\mathbf{x}_{hr},\mathbf{x}_{t_{i}})+\beta\omega_{ht_{i}}\end{aligned}$
@@
where $\gamma$ is an additive margin, a temperature parameter $\tau$ adjusts the importance of negatives, a structural hardness factor $\omega_{ht_{i}}$ stands for how hard a negative triple ($h,r,t_{i}$) is in terms of the structural relation between $h$ and $t_{i}$ in $\mathcal{G}$, and $\beta$ is a trainable parameter that adjusts the relative importance of $\omega_{ht_{i}}$. We define $\omega_{ht_{i}}$ as the reciprocal of the distance (i.e., length of the shortest path) between $h$ and $t_{i}$ to impose a larger $\omega_{ht_{i}}$ to the negative triple with a shorter distance between $h$ and $t_{i}$ in $\mathcal{G}$, serving as a hard negative triple.
@@
Since computing the exact distance between every head and every tail in $\mathcal{B}$ may spend considerable time, we calculate the approximate distance between $h$ and $t_{i}$, i.e., the multiplication of two distances: (d1) the distance between $h$ and head $h_{c}$ of the center triple of $\mathcal{B}$, and (d2) the distance between $t$ and $h_{c}$. Thus, the distance from $h_{c}$ to every entity in $\mathcal{B}$ is pre-computed before training, the multiplication between the two distances is performed in parallel during training, requiring a minimal computational overhead. [^6] We conducted experiments using the exact distance and different approximate distances, e.g., the sum of (d1) and (d2), but the performance gap between all the methods is very marginal.
4.4. Frequency-aware Mini-batch Training
@@
Furthermore, the frequency of every entity that occurs in a set $\mathcal{T}$ of triples varies; this frequency distribution follows the power law. However, the text-based methods have shown difficulties in addressing the discrepancy between this long-tailed distribution of the KG and the distribution of a mini-batch. Figure 5 shows the frequency distributions of the original KG, 100 randomly-sampled mini-batches from $\mathcal{T}$, and those from SaaM, where $y$-axis denotes the frequency ratio of entity occurrences, and entities in $x$-axis are sorted in the ascending order of their degrees in KG. Given $h$ and $t$, in fact, KGC can be regarded as the multi-class classification task to predict a correct tail $t$. From this perspective, severe class imbalance originating from the long-tail distribution (red lines) in the KG may result in the less-skewed distribution (green lines) for 100 randomly-sampled mini-batches from $\mathcal{T}$, due to a limited number of entities in the mini-batches. In addition, we observe the nearly uniform distribution for 100 mini-batches selected by SaaM (blue lines), since SaaM is likely to give preference to low-degree entities. Therefore, we apply the importance sampling scheme to associate the expected error of $\mathcal{L}_{(}h,r,t)$ with the long-tailed original domain of the KG:
$\mathbb{E}_{p_{o}{(h,r,t)}}\left[\mathcal{L}_{(}h,r,t)\right]$ $=\mathbb{E}_{p_{s}{(h,r,t)}}\left[\frac{p_{o}(h,r,t)}{p_{s}(h,r,t)}\mathcal{L}_{(}h,r,t)\right]$
$=\mathbb{E}_{p_{s}{(h,r,t)}}\left[\frac{p_{o}(t)p(h,r|t)}{p_{s}(t)p(h,r|t)}\mathcal{L}_{(}h,r,t)\right]$
$:=\mathbb{E}_{p_{s}{(h,r,t)}}\left[\psi_{\mathcal{G}}(t)\mathcal{L}_{(h,r,t)}\right]$
Figure 5. Frequency distributions of entities for original KG, 100 mini-batches randomly sampled from $\mathcal{T}$, and those randomly sampled by SaaM on WN18RR and FB15k-237. The entities are sorted in the ascending order of their degrees.
where $p_{s}(h,r,t)$ and $p_{o}(h,r,t)$ are probabilities of sampling $(h,r,t)$ by SaaM and randomly in $\mathcal{T}$ respectively. A weighting factor $\psi_{\mathcal{G}}(t)$ of tail $t$ denotes $p_{o}(t)/p_{s}(t)$. Assume the degree-proportional $p_{o}(t)=|N_{t}|/2|\mathcal{T}|$ and the nearly-uniform distribution $p_{s}(t)=1/|\mathcal{E}|$ where $|\mathcal{T}|=d_{avg}|\mathcal{E}|/2$ with $d_{avg}$ being the average number of neighbors for every entity. Consequently, $p_{o}(t)/p_{s}(t)=|N_{t}|/d_{avg}\propto|N_{t}|$.
To encourage PLMs to more sensitively adapt to varying relations for many triples with the identical head and to eliminate the discrepancy of the two distributions, we propose to reweight the importance of each sample loss via $\psi_{\mathcal{G}}(t)$ above, so the frequency distribution of $\mathcal{B}$ becomes close to that of $\mathcal{T}$. For each triple in a mini-batch $\mathcal{B}$ in SaaM, mini-batch loss $\mathcal{L}_{\mathcal{B}}$ is defined as:
$\mathcal{L}_{\mathcal{B}}=\sum_{(h,r,t)\in\mathcal{B}}\psi_{\mathcal{G}}(t)\mathcal{L}_{(h,r,t)}$
where $\psi_{\mathcal{G}}(t)$ is defined as $log(|N_{t}|+1)$ where $N_{t}$ is a set of $t$’s neighbors in the KG. Applying logarithm can prevent $\psi_{\mathcal{G}}(t)$ from growing extremely large for a few entities with a huge $|N_{t}|$, and adding one can prevent $\psi_{\mathcal{G}}(t)$ from becoming zero when $|N_{t}|=1$. To sum up, $\psi_{\mathcal{G}}(t)$ ensures that the triples with larger-degree tails contribute more significantly to $\mathcal{L}_{\mathcal{B}}$.
5. EXPERIMENTS
5.1. Experimental Setup
For evaluation we adopt widely-used KG datasets WN18RR, FB15k-237 and Wikidata5M. Table 1 shows their statistics, and more details are provided in Appendix C.
For every incomplete triple in the test set, we compute mean reciprocal rank (MRR) and Hits@$k$ where $k\in\{1,3,10\}$ as evaluation metrics based on the rank of the correct entity among all the entities in the KG. We use the mean of the forward and backward prediction results as the final performance measure. The hyperparameters of SATKGC are set based on the experimental results in Appendix F. Further implementation details are described in Appendix E.
@@ 5.1. Experimental Setup
Table 1. Statistics of datasets.
Approach WN18RR FB15k-237
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
@@
TuckER (Balazevic et al., 2019) 0.466 0.432 0.478 0.518 0.361 0.265 0.391 0.538
RotatE (Sun et al., 2019) 0.471 0.421 0.490 0.568 0.335 0.243 0.374 0.529
KGTuner (Zhang et al., 2022) 0.481 0.438 0.499 0.556 0.345 0.252 0.381 0.534
KG-Mixup (Shomer et al., 2023) 0.488 0.443 0.505 0.541 0.359 0.265 0.395 0.547
[…]
CompoundE (Ge et al., 2023) 0.492 0.452 0.510 0.570 0.350 0.262 0.390 0.547
KRACL (Tan et al., 2023) 0.529 0.480 0.539 0.621 0.360 0.261 0.393 0.548
CSProm-KG (Chen et al., 2023) 0.569 0.520 0.590 0.675 0.355 0.261 0.389 0.531
@@
StAR (Wang et al., 2021b) 0.398 0.238 0.487 0.698 0.288 0.195 0.313 0.480
HaSa (Zhang et al., 2023c) 0.535 0.446 0.587 0.711 0.301 0.218 0.325 0.482
KG-S2S (Chen et al., 2022) 0.572 0.529 0.595 0.663 0.337 0.255 0.374 0.496
SimKGC (Wang et al., 2022) 0.671 0.580 0.729 0.811 0.340 0.252 0.365 0.515
[…]
SATKGC (w/o SaaM) 0.673 0.595 0.728 0.813 0.349 0.256 0.367 0.520
SATKGC (w/o PCL, FMT) 0.676 0.608 0.722 0.820 0.355 0.261 0.386 0.537
SATKGC (w/o FMT) 0.680 0.611 0.729 0.823 0.351 0.265 0.389 0.541
SATKGC (w/o PCL) 0.686 0.623 0.740 0.827 0.366 0.272 0.396 0.545
SATKGC 0.694 0.626 0.743 0.833 0.370 0.279 0.403 0.550
Table 2. KGC results for the WN18RR, and FB15k-237 datasets. “PCL” and “FMT” refer to Proximity-aware Contrastive Learning and Frequency-aware Mini-batch training respectively. The best and second-best performances are denoted in bold and underlined respectively. $\dagger$: numbers are from Liu et al. (2024) $\ddagger$: numbers are from Qiao et al. (2023).
5.2. Main Results
We compare SATKGC with existing embedding-based and text-based approaches. Table 2 shows the results on WN18RR, and FB15k-237, and Table 3 shows the results on Wikidata. Due to the page limit, we present comparison with recent models in these two tables. Additional comparison results can be found in Appendices N and M.
SATKGC denotes the bi-encoder architectures trained by our learning framework in Figure 3. SATKGC consistently outperforms all the existing methods on all the datasets. SATKGC demonstrates significantly higher MRR and Hits@1 than other baselines, with Hits@1 improving by 5.03% on WN18RR and 5.28% on FB15k-237 compared to the existing state-of-the-art models.
@@ 5.2. Main Results
TransE (Bordes et al., 2013) 0.253 0.170 0.311 0.392 - - - -
TuckER (Balazevic et al., 2019) 0.285 0.241 0.314 0.377 - - - -
RotatE (Sun et al., 2019) 0.290 0.234 0.322 0.390 - - - -
KGTuner (Zhang et al., 2022) 0.305 0.243 0.331 0.397 - - - -
@@
DKRL (Xie et al., 2016) 0.160 0.120 0.181 0.229 0.231 0.059 0.320 0.546
KEPLER (Wang et al., 2021a) 0.212 0.175 0.221 0.276 0.403 0.225 0.517 0.725
BLP-ComplEx (Daza et al., 2021) - - - - 0.491 0.261 0.670 0.881
BLP-SimplE (Daza et al., 2021) - - - - 0.490 0.283 0.641 0.868
SimKGC (Wang et al., 2022) 0.358 0.313 0.376 0.441 0.714 0.609 0.785 0.917
SATKGC 0.408 0.366 0.425 0.479 0.767 0.675 0.815 0.931
Table 3. KGC results for Wikidata5M-Trans (transductive setting) and Wikidata5M-Ind (inductive setting). The results for embedding-based approach on Wikidata5M-ind are missing as they cannot be used in the inductive setting. Additionally, BLP-ComplEx (Daza et al., 2021) and BLP-SimplE (Daza et al., 2021) results on Wikidata5M-Trans are missing because they are inherently targeted for inductive KGC.
As shown in Table 3, SATKGC demonstrates its applicability to large-scale KGs, and achieves strong performance in both inductive and transductive settings. [^7] Baselines listed in Table 2 but not in Table 3 could not be evaluated on Wikidata5M due to out-of-memory for StAR and CSProm-KG, or unreasonably large training time, i.e., more than 100 hours expected, for the remaining baselines. SATKGC shows an improvement of 7.42% in MRR and 10.84% in Hits@1 compared to the existing state-of-the-art model on Wikidata5M-Ind. On Wikidata5M-Trans, SATKGC achieves an improvement of 13.97% in MRR and 16.93% in Hits@1 over the previous best-performing model. [^8] Wikidata5M-Ind shows better performance than Wikidata5M-Trans, because a model ranks 7,475 entities in the test set for Wikidata5M-Ind while ranking about 4.6 million entities for Wikidata5M-Trans. In the transductive setting, performance degrades in the order of WN18RR, Wikidata5M-Trans, and FB15k-237, showing that a higher average degree of entities tends to negatively affect performance. A more detailed analysis of performance differences across the dataset is described in Appendix I.
@@
KoPA (Zhang et al., 2023a) 0.483 0.344 0.559 0.721
SATKGC 0.569 0.450 0.637 0.792
@@
Table 4. KGC results on FB15K-237N. A correct tail is ranked among 1,000 randomly-selected entities due to the long inference time of KoPA.
@@
To compare our framework employing MPNets with a LLM-based model, we evaluate the KGC performance of KoPA (Zhang et al., 2023a), which adopts Alpaca (Taori et al., 2023) fine-tuned with LoRA (Hu et al., 2021) as its backbone, on the FB15k-237N dataset (Zhang et al., 2023a). [^9] We adopt FB15k-237N in Table 4 to compare our model with KoPA, because the official implementation of KoPA provides the pretrained model parameters for not FB15k-237 but FB15K-237N. Since KoPA cannot perform inference for all queries within a reasonable time, i.e., 111 hours expected, we rank the correct tail among 1,000 randomly selected entities for both KoPA and SATKGC. As shown in Table 4, SATKGC outperforms KoPA on all metrics, demonstrating that LLMs do not necessarily produce superior results on KGC.
5.3. Ablation Study
To demonstrate the contribution of each component of our method, we compare the results across five different settings, including SATKGC, as shown in Table 2. In SATKGC (w/o SaaM), instead of using SaaM, a mini-batch of triples is randomly selected without replacement, while both Proximity-aware Contrastive Learning and Frequency-aware Mini-batch training are applied. SATKGC (w/o PCL, FMT) refers to applying only SaaM and using the original InfoNCE loss. SATKGC (w/o FMT) applies both SaaM and Proximity-aware Contrastive Learning, and SATKGC (w/o PCL) applies both SaaM and Frequency-aware Mini-batch Training. The results show that SATKGC (w/o SaaM) performs worse than SATKGC, indicating that substituting SaaM with random sampling significantly hurts the performance. SATKGC (w/o PCL, FMT) already achieves higher Hits@1 than other baselines on WN18RR, highlighting that SaaM alone leads to performance improvement [^10] Performance improvements are also observed when we apply SaaM to embedding-based RotatE and text-based StAR, as described in Appendix G.. Between SATKGC (w/o PCL) and SATKGC (w/o FMT), SATKGC (w/o PCL) achieves higher performance, indicating that Frequency-aware Mini-batch Training contributes more than Proximity-aware Contrastive Learning.